# Optimizing energy conversion with nonthermal resources in steady-state quantum devices

### v1

@@ I Introduction
Conventional thermoelectric heat engines and refrigerators are energy converters that operate between two (or more) thermal reservoirs of different temperatures and electrochemical potentials. Thermoelectric energy conversion in nanoscale and mesoscopic conductors has recently attracted significant attention [1, 2, 3, 4, 5, 6, 7], driven by its versatile potential for efficient power production and cooling [8]. Due to the controllable energy-filtering properties of the conductor, nanoelectronic devices can act as highly tunable and efficient thermoelectric energy harvesters [9, 10, 11]. This has also been demonstrated in recent experiments using quantum dots or molecular junctions [12, 13, 14, 15, 16, 17].
In these small-scale devices, mostly operating at low temperatures, the thermalization length can however easily exceed the system size [4]. The electronic distributions in the contacts are hence possibly nonthermal. This means that they cannot be characterized by temperatures and electrochemical potentials. Nonthermal electronic distributions can arise from the interplay between competing environments [18, 19], from irradiation by light fields [20, 21, 22, 23] or even due to an external driving of some nearby operated device. They can also be created by controlled mixing of different thermal distributions [24, 25, 26] or by correlations between system and bath due to their strong coupling [27]. In fact, nonthermal distributions can be experimentally measured and fully characterized, for example by state spectroscopy [28]. Nonthermal distributions can be exploited as resource for energy conversion processes, potentially offering even greater advantages than their thermal counterparts [29, 30, 31, 32]. However, bounds on the optimal performance of energy conversion processes mostly resort to the thermal properties of the resources, such as the fundamental Carnot bound for the efficiency or the Curzon-Ahlborn bound [5]. There already exist some guidelines for optimizing the performance of nanoscale energy conversion processes in nonequilibrium devices—primarily targeting coupling to macroscopic thermal reservoirs [33, 11, 34, 35, 36]. However, corresponding directives remain lacking for energy conversion processes that specifically exploit the unique features of nonthermal resources.
@@
We address this challenge by optimizing the performance of coherent electronic conductors with two contacts, where one of the contacts is characterized by an arbitrary nonthermal distribution, see panel(b) of Fig. 1. Scattering theory is used to describe the coherent charge transmission between the reservoirs, which is experimentally achievable at the nanoscale [37, 38, 39, 16, 40, 41, 42]. This means that our results are applicable to any type of coherent conductor where many-body interactions are weak enough that they can be treated within a mean-field approach and where all inelastic processes that cause dissipation or decoherence occur in the reservoirs [43]. To quantify the performance of this generic nonthermal situation, we use general transport currents to define the input as well as the output of the conversion process and the resulting generalized efficiencies, since heat flows are often not a meaningful concept any longer. As input current from the resource, entropy currents are a meaningful quantity; a resource is “used up” when its entropy is increased, allowing by virtue of the second law that the entropy is decreased elsewhere in the device.
In addition to generic output currents and the related efficiencies, we also optimize the precision of these generic output currents as well as trade-off relations between precision and efficiency in the spirit of recently studied thermodynamic uncertainty relation [44, 45, 46].
Concretely, our optimization aims at finding the ideal transmission properties of the conductor with respect to the mentioned performance quantifiers, provided a fixed desired output current. To treat this constraint of a given output current, we use a Lagrange multiplier method, extending the strategy of Refs. [33, 11, 35]. This further allows us to expand our optimization to additional tunable parameters, such as externally applied gate potentials.
We find that the performance of the device is optimized by a series of boxcar-shaped energy filters. We identify crossing points between characteristic spectral quantities which set the positions and widths of these energy filters. This provides guidelines for the design of nanoelectronic devices for energy conversion processes with nonthermal resources. By working out the ideal performance, we also provide bounds for realistic devices. Consequently, this means that we identify the largest output current, the highest efficiency at given output current and the best precision that one can possibly achieve in a two-terminal coherent conductor, exploiting a given nonthermal resource.
@@ II Model and Theoretical Approach
This section introduces the considered model for the two-terminal coherent conductor with a nonthermal resource. Furthermore, all relevant performance quantifiers are defined and we show how to evaluate using scattering theory for quantum transport.
@@ II.1 Two-terminal conductor with a nonthermal contact
The subject of this study is a nanoelectronic conductor operating coherently between two terminals, as sketched in Fig. 1. The two macroscopic contacts have very different characteristics. The left contact is described by a thermal distribution, where the occupation probability of particles at energy $\varepsilon$ is given by a Fermi function $f(\varepsilon)=\left(1+\exp[(\varepsilon-\mu)/k_{\mathrm{B}}T]\right)^{-1}$, with temperature $T$ and electrochemical potential $\mu$ (set to reference throughout the paper as $\mu=0$). By contrast, the distribution of the right contact can be nonthermal (see panel (b) compared to panel (a)). This means that the particles’ occupation probabilities $g(\varepsilon)$, are generally not given by a Fermi function, but only need to fulfill the constraint $0\leq g(\varepsilon)\leq 1$, due to the Pauli exclusion principle. Throughout this paper, we assume that both electron distribution are given and constant, since the device is steady-state.
The nonthermal contact is considered to be the resource for energy conversion in this manuscript (and compared to an equivalent hot thermal contact). Such steady-state nonthermal distributions occur when the length scale of the device is smaller than its thermalization length [4] or the extraction times of particles are much shorter than the thermalization time. This happens for example in hot carrier solar cells [18, 47, 19] or other devices where different environments compete or are mixed with each other [24, 29, 28] or in systems steadily irradiated by coherent light [20, 21, 22, 23, 48], but they can also be an effect of energy filtering [49, 50]. Characteristic examples are discussed in Sec. IV.
To utilize the resource to generate a useful output, the energy-dependent transmission properties of the central conductor are exploited. Many-body interactions in the coherent central conductor are assumed to be weak [43]. Therefore, it is possible to capture the properties of the conductor by a transmission probability, $\mathcal{D}(\varepsilon)\in[0,1]$. In Fig. 1, the conductor acts as an energy filter where $\mathcal{D}(\varepsilon)=1$ where electrons are transmitted and $\mathcal{D}(\varepsilon)=0$ where they are reflected.
@@
The transmission function is the quantity that is optimized in Sec. III. All other quantities and functions are assumed to be given, with the exception of a possible second variable in Sec. III.5. This perspective reflects designing a nanoelectronic conductor for a particular task and known, constant electronic distributions.
@@ II.2 Transport observables from scattering theory
The device of Fig. 1 acts as an energy converter where energy is carried by electrons, or holes in the Fermi sea. Heat transport due to phonons can typically be neglected in low-temperature experiments [4]. Note however that the nonthermal distributions treated in this manuscript can indeed be the result of interactions between electrons and phonons or photons.
To analyze the performance of the two-terminal conductor as an energy converter, both the output of the conversion and the input provided by the nonthermal resource need to be characterized. Therefore, we calculate the particle, energy, and entropy currents out of a contact flowing towards the scatterer (the coherent conductor). Assuming that interaction effects beyond the mean-field level can be neglected, scattering theory is used to calculate these transport quantities, given by [51]
@@
Here, $z(\varepsilon)$ is the current coefficient which describes the type of current in question. Some notable examples are $z(\varepsilon)=\pm 1/h$ for the particle current $I^{N,\mathrm{L/R}}$ out of contact L/R and $z(\varepsilon)=\pm(\varepsilon/h)$ for the energy current $I^{\varepsilon,\mathrm{L/R}}$ out of contact L/R. Further, $z(\varepsilon)=(k_{\mathrm{B}}/h)\log[g(\varepsilon)/(1-g(\varepsilon)]$ gives the entropy current $I^{\Sigma,\mathrm{R}}$ out of the right contact and $z(\varepsilon)=-(k_{\mathrm{B}}/h)\log[f(\varepsilon)/(1-f(\varepsilon))]$ the entropy current $I^{\Sigma,\mathrm{L}}$ out of the left contact [32, 52].
These currents fulfill the thermodynamic laws [4], which implies the following conservation laws or limits:
$I^{\varepsilon,\mathrm{L}}+I^{\varepsilon,\mathrm{R}}=0$ (2a)
$I^{N,\mathrm{L}}+I^{N,\mathrm{R}}=0$ (2b)
$-I^{\Sigma,\mathrm{L}}-I^{\Sigma,\mathrm{R}}\geq 0$ (2c)
Importantly, particle and energy currents are conserved and therefore just change sign when considered in the different contacts. This is different for the entropy current, where $k_{\mathrm{B}}\log[\frac{g(\varepsilon)}{1-g(\varepsilon)}]$ and $k_{\mathrm{B}}\log[\frac{f(\varepsilon)}{1-f(\varepsilon)}]$ can be different, or in other words, entropy can be produced, only being constrained from below by the second law. Since the left contact is always thermal, we furthermore have $k_{\mathrm{B}}\log[\frac{f(\varepsilon)}{1-f(\varepsilon)}]=(\varepsilon-\mu)/T$. This means that the entropy current is directly proportional to the heat current, $TI^{\Sigma,\mathrm{L}}=I^{Q,\mathrm{L}}$, which is obtained [51, 4] by setting $z(\varepsilon)=(\varepsilon-\mu)/T$ in Eq. (1). By contrast, in the nonthermal contact the heat current is generally not well-defined, due to the absence of a consistent definition of temperature and electrochemical potential.
The average current, given in Eq. (1), can be decomposed as an energy integral over the spectral currents. The spectral currents are given by
@@
Since the precision is a further performance quantifier of interest, in addition to the expectation values of the currents, their fluctuations will be considered. The zero-frequency noise of a current $I^{z}$ obtained from scattering theory has the explicit expression [53, 52]
$S^{z}$ $=$ $\int_{-\infty}^{\infty}d\varepsilon\,z^{2}(\varepsilon)\mathcal{D}(\varepsilon)(g(\varepsilon)g^{-}(\varepsilon))+f(\varepsilon)f^{-}(\varepsilon))$ (5)
@@
$s^{z}(\varepsilon)\equiv s^{z}_{\mathrm{thermal}}(\varepsilon)=z^{2}(\varepsilon)\mathcal{D}(\varepsilon)(g(\varepsilon)g^{-}(\varepsilon)+f(\varepsilon)f^{-}(\varepsilon)),$ (6)
@@ II.3 Performance quantifiers
Our objective is to optimize the performance of energy converters. Therefore, a crucial first step is to identify meaningful performance quantifiers. The first important aspect for the performance is naturally the actual desired output. This is quantified here as a generic current, $I^{x}$, see the definition in Eq. (1), with current factor $x(\varepsilon)$. For the matter of clarity and without loss of generality, we here always take the output current to be positive, $I^{x}\geq 0$. Consequently optimization of the output current, which we present in Sec. III.1 always consists of maximizing $I^{x}$. Concretely, this means that if the desired output is negative following the definition in Eq. (1), one needs to set $x(\varepsilon)\to-z(\varepsilon)$ for the optimization procedure described below. This ensures that the total and spectral currents convey the same physical meaning.
@@
The strategy of how to define a meaningful, dimensionless efficiency becomes clear from the concrete example considered in Sec. IV, where we study the cooling power in the left contact. Since the right, resource contact is not thermal, the resource is typically not power or a heat current, as one would expect from a standard Peltier element or absorption refrigerator. Instead, motivated by the constraints put on energy conversion by the second law, we aim at characterizing the resource by an entropy current $I^{y}\rightarrow-I^{\Sigma,\mathrm{R}}$ (which is positive when entropy is produced in contact R). The constraints from the first law ensure energy and particle conservation, but this cannot yield insight into the heat transfer process. In the thermal (left) contact, the entropy reduction is proportional to the desired cooling power, namely to the heat current leaving the reservoir $I^{x}\rightarrow I^{Q,\mathrm{L}}=TI^{\Sigma,\mathrm{L}}$. The increase in entropy, $I^{\Sigma,\mathrm{R}}>0$, implies that the right side acts as a resource (namely allowing by means of the second law that entropy is reduced elsewhere). Using the constraint by the second law (2c), we hence have $\eta=\frac{1}{T}\frac{I^{Q,\mathrm{L}}}{-I^{\Sigma,\mathrm{R}}}\leq 1$ with $c_{\eta}=1/T$. Alternatives for quantifying the resource that have been considered previously could be the free-energy current [30, 54, 19] (with respect to the temperature of the thermal contact), $I^{y}\rightarrow I^{\varepsilon,\mathrm{R}}-TI^{\Sigma,\mathrm{R}}$.
@@
Furthermore, it is important in nanoscale conductors that the output of the energy-conversion process is precise. We hence aim at reducing the noise, where the performance quantifier is the noise of the current $I^{x}$ itself, $S^{x}$, which needs to be reduced, compared to the actual desired (average) output. This can be quantified in the noise-to-signal ratio
$\mathrm{NSR}=c_{\mathrm{NSR}}\frac{S^{x}}{I^{x}}\ .$ (8)
@@
where $c_{\mathrm{NSR}}$ is again a normalization factor for the dimensionality. For charge currents, this ratio is known as the Fano factor [53], where the constant takes the value $c_{\mathrm{NSR}}=1/2e$.
@@
It is well known from the output and the efficiency of standard engines that not both of these goals can be reached simultaneously, but that trade-offs need to be found. Here, we will consider one particular trade-off inspired by the thermodynamic uncertainty relation (TUR) [44, 46], stating that precision is bounded by entropy production. We will therefore also consider the TUR-inspired factor
$A^{x,y}=c_{A}\frac{S^{x}I^{y}}{(I^{x})^{2}}$ (9)
as a performance quantifier, where we again introduce a normalization factor $c_{A}$. For comparison, the TUR is expressed as
$A_{\mathrm{TUR}}^{x,\sigma}=\frac{S^{x}\sigma}{(I^{x})^{2}},$ (10)
with $\sigma=-I^{\Sigma,L}-I^{\Sigma,R}$ the total entropy production rate. Our definition is a more general trade-off relation, where $I^{y}$ is not specified from the start. For classical processes the TUR-inspired factor given in Eq. (9) equals (10) and is limited by $2k_{B}$, when considering the full entropy production, $I^{y}\to-I^{\Sigma,\mathrm{L}}-I^{\Sigma,\mathrm{R}}$ and setting $c_{A}$ to 1.

### v2

@@ I Introduction
Conventional thermoelectric heat engines and refrigerators are energy converters that operate between two (or more) thermal reservoirs of different temperatures and electrochemical potentials. Thermoelectric energy conversion in nanoscale and mesoscopic conductors has recently attracted significant attention [Sothmann_2013, Haupt2013, QTDBook2018, Benenti2017Jun, Whitney2018May, Kirchberg2022, Campbell2025Apr], driven by its versatile potential for efficient power production and cooling [Giazotto2006Mar]. Due to the controllable energy-filtering properties of the conductor, nanoelectronic devices can act as highly tunable and efficient thermoelectric energy harvesters [Hicks1993May, Mahan1996Jul, Whitney2015]. This has also been demonstrated in recent experiments using quantum dots or molecular junctions [Reddy2007, Josefsson2018Oct, Thierschmann2015Oct, Roche2015Apr, Jaliel2019Sep, Gehring2021Apr].
In these small-scale devices, mostly operating at low temperatures, the thermalization length can however easily exceed the system size [Benenti2017Jun]. The electronic distributions in the contacts are hence possibly nonthermal. This means that they cannot be characterized by temperatures and electrochemical potentials. Nonthermal electronic distributions can arise from the interplay between competing environments [Konig2010Sep, Tesser2023], from irradiation by light fields [Lloyd-Hughes2021Jul, Maiuri2020Jan, Uehlein2025Mar, Cherubim2019] or even due to an external driving of some nearby operated device. They can also be created by controlled mixing of different thermal distributions [Altimiras2010Nov, Kovrizhin2012Sep, leSueur2010Jul] or by correlations between system and bath due to their strong coupling [Aguilar2024], and even in other types of non-electronic systems, such as in trapped ions [Rossnagel2014] or active matter like bacteria  [Krishnamurthy2016Dec, Wiese2024]. In electronic systems, nonthermal distributions can in fact be experimentally measured and fully characterized, for example by state spectroscopy [Altimiras2010Oct]. Nonthermal distributions can be exploited as resource for energy conversion processes, potentially offering even greater advantages than their thermal counterparts [Sanchez2019Nov, Hajiloo2020Oct, Ciliberto2020Nov, Deghi2020Jul]. However, bounds on the optimal performance of energy conversion processes mostly resort to the thermal properties of the resources, such as the fundamental Carnot bound for the efficiency or the Curzon-Ahlborn bound [Whitney2018May]. There already exist some guidelines for optimizing the performance of nanoscale energy conversion processes in nonequilibrium devices—primarily targeting coupling to macroscopic thermal reservoirs [Whitney2014Apr, Whitney2015, Landi2025, Chrirou2025Jul, Ryu_2022]. However, corresponding directives remain lacking for energy conversion processes that specifically exploit the unique features of nonthermal resources.
@@
We address this challenge by optimizing the performance of coherent electronic conductors with two contacts, where one of the contacts is characterized by an arbitrary nonthermal distribution, see panel (b) of Fig. 1, which is maintained like this in steady state. Scattering theory is used to describe the coherent charge transmission between the contacts, which is experimentally achievable at the nanoscale [VanHouten1992, Brantut2012, Brantut2013, Jaliel2019Sep, Putzke2020, Kobayashi2021Sep, Balduque2025Apr]. This means that our results are applicable to any type of coherent conductor where many-body interactions are weak enough that they can be treated within a mean-field approach and where all inelastic processes that cause dissipation or decoherence occur in the contacts [Christen_1996].
To quantify the performance of this generic nonthermal situation, we use general transport currents to define the input as well as the output of the conversion process and the resulting generalized efficiencies, since heat flows are often not a meaningful concept any longer. In fact, defining work in quantum systems interacting with different types of environments is often challenging [Talkner2020, Landi2021Sep, Silva2021Oct]. Specifically in quantum transport, work and heat are defined in terms of electrochemical potentials [Benenti2017Jun], which are not defined for nonthermal distributions. A current that can instead be defined is the entropy current [Deghi2020Jul, Acciai2024Feb, Tesser2025], the change of entropy in a terminal, which is found by accessing the electronic distributions of the terminals. As the input current from the resource, the entropy current is a meaningful quantity; a resource is “used up” when its entropy is increased, allowing by virtue of the second law that the entropy is decreased elsewhere in the device.
In addition to generic output currents and the related efficiencies, we also optimize the precision of these generic output currents as well as trade-off relations between precision and efficiency in the spirit of recently studied thermodynamic uncertainty relations [Barato2015Apr, Gingrich2016, Pietzonka2018May].
Concretely, our optimization aims at finding the ideal transmission properties of the conductor with respect to the mentioned performance quantifiers, provided a fixed desired output current. To treat this constraint of a given output current, we use a Lagrange multiplier method, extending the strategy of Refs. [Whitney2014Apr, Whitney2015, Chrirou2025Jul]. This further allows us to expand our optimization to additional tunable parameters, such as externally applied gate potentials.
We find that the performance of the device is optimized by a series of boxcar-shaped energy filters. We identify crossing points between characteristic spectral quantities which set the positions and widths of these energy filters. This provides guidelines for the design of nanoelectronic devices for energy conversion processes with nonthermal resources. By working out the ideal performance, we also provide bounds for realistic devices. Consequently, this means that we identify the largest output current, the highest efficiency at given output current and the best precision that one can possibly achieve in a two-terminal coherent conductor exploiting a given nonthermal resource.
@@ II Model and Theoretical Approach
This section introduces the considered model for the two-terminal coherent conductor with a nonthermal resource. Furthermore, all relevant performance quantifiers are defined and we show how to evaluate them using scattering theory for quantum transport.
@@ II.1 Two-terminal conductor with a nonthermal contact
The subject of this study is a nanoelectronic conductor operating coherently between two terminals, as sketched in Fig. 1. The two macroscopic contacts have very different characteristics. The left contact is described by a thermal distribution, where the occupation probability of particles at energy $\varepsilon$ is given by a Fermi function $f(\varepsilon)=\left(1+\exp[(\varepsilon-\mu)/k_{\mathrm{B}}T]\right)^{-1}$, with temperature $T$ and electrochemical potential $\mu$ (set to reference throughout the paper as $\mu=0$). By contrast, the distribution of the right contact can be nonthermal (see panel (b) compared to panel (a)). This means that the particles’ occupation probabilities $g(\varepsilon)$ are generally not given by a Fermi function, but only need to fulfill the constraint $0\leq g(\varepsilon)\leq 1$, due to the Pauli exclusion principle. Throughout this paper, we assume that both electron distributions are given and constant, since the device is kept in a steady-state situation.
The nonthermal contact is considered to be the resource for energy conversion in this manuscript (and compared to an equivalent hot thermal contact). Such steady-state nonthermal distributions occur when the length scale of the device is smaller than its thermalization length [Benenti2017Jun] or the extraction times of particles are much shorter than the thermalization time. This happens for example in hot carrier solar cells [Konig2010Sep, Ahmed_2021, Tesser2023] or other devices where different environments compete or are mixed with each other [Altimiras2010Nov, Sanchez2019Nov, Altimiras2010Oct] or in systems steadily irradiated by coherent light [Lloyd-Hughes2021Jul, Maiuri2020Jan, Uehlein2025Mar, Cherubim2019, Sarkar_2022], but nonthermal distributions can also be a result of energy filtering [Heremans2008Jul, Bahk2013Feb]. The act of energy filtering by the conductor, which we here exploit for cooling, can therefore in principle also affect the distributions in the terminals, which can be taken into account in self-consistent manner [Tesser2023, Bertin-Johannet2025Jul]. However, accounting for this would require knowing the specific circumstances under which the nonthermal distribution is created; here, we instead focus on generic nonthermal and thermal distributions in the terminals. Characteristic examples of such nonthermal distributions are discussed in Sec. IV.
To utilize the resource to generate a useful output, the energy-dependent transmission properties of the central conductor are exploited. Many-body interactions in the coherent central conductor are assumed to be weak [Christen_1996]. Therefore, it is possible to capture the properties of the conductor by a transmission probability, $\mathcal{D}(\varepsilon)\in[0,1]$. In Fig. 1, the conductor acts as an energy filter with $\mathcal{D}(\varepsilon)=1$ where electrons are transmitted and $\mathcal{D}(\varepsilon)=0$ where they are reflected.
@@
The transmission function is the quantity that is optimized in Sec. III. All other quantities and functions are assumed to be given, with the exception of a possible second variable in Sec. III.5. This perspective reflects designing a nanoelectronic conductor for a particular task and known constant electronic distributions.
@@ II.2 Transport observables from scattering theory
The device in Fig. 1 acts as an energy converter where energy is carried by electrons or holes in the Fermi sea. Heat transport due to phonons can typically be neglected in low-temperature experiments [Benenti2017Jun]. Note however that the nonthermal distributions treated in this manuscript could indeed be the result of interactions between electrons and phonons or photons.
To analyze the performance of the two-terminal conductor as an energy converter, both the output of the conversion and the input provided by the nonthermal resource need to be characterized. Therefore, we calculate the particle, energy, and entropy currents out of a contact flowing towards the scatterer (the coherent conductor). Assuming that interaction effects beyond the mean-field level can be neglected, scattering theory is used to calculate these transport quantities. The average current is derived from a current operator $\hat{I}^{z}(t)$ by taking the expectation value, yielding [Moskalets2011Sep]
@@
Here, $z(\varepsilon)$ is the current coefficient which describes the type of current in question. Some notable examples are $z(\varepsilon)=\pm 1/h$ for the particle current $I^{N,\mathrm{L/R}}$ out of contact L/R and $z(\varepsilon)=\pm(\varepsilon/h)$ for the energy current $I^{\varepsilon,\mathrm{L/R}}$ out of contact L/R. Further, $z(\varepsilon)=(k_{\mathrm{B}}/h)\ln[g(\varepsilon)/(1-g(\varepsilon)]$ gives the entropy current $I^{\Sigma,\mathrm{R}}$ out of the right contact and $z(\varepsilon)=-(k_{\mathrm{B}}/h)\ln[f(\varepsilon)/(1-f(\varepsilon))]$ the entropy current $I^{\Sigma,\mathrm{L}}$ out of the left contact [Deghi2020Jul, Acciai2024Feb].
These currents fulfill the thermodynamic laws [Benenti2017Jun, Topp2015, Nenciu2007], which implies the following conservation laws or limits:
$I^{\varepsilon,\mathrm{L}}+I^{\varepsilon,\mathrm{R}}=0,$ (2a)
$I^{N,\mathrm{L}}+I^{N,\mathrm{R}}=0,$ (2b)
$-I^{\Sigma,\mathrm{L}}-I^{\Sigma,\mathrm{R}}\geq 0.$ (2c)
Importantly, particle and energy currents are conserved and therefore just change sign when considered in the different contacts. This is different for the entropy current, where $k_{\mathrm{B}}\ln[\frac{g(\varepsilon)}{1-g(\varepsilon)}]$ and $k_{\mathrm{B}}\ln[\frac{f(\varepsilon)}{1-f(\varepsilon)}]$ can be different, or in other words, entropy can be produced, only being constrained from below by the second law. Since the left contact is always thermal, we furthermore have $k_{\mathrm{B}}\ln[\frac{f(\varepsilon)}{1-f(\varepsilon)}]=(\varepsilon-\mu)/T$. This means that the entropy current is directly proportional to the heat current, $TI^{\Sigma,\mathrm{L}}=I^{Q,\mathrm{L}}$, which is obtained [Moskalets2011Sep, Benenti2017Jun] by setting $z(\varepsilon)=(\varepsilon-\mu)/T$ in Eq. (1). By contrast, in the nonthermal contact the heat current is generally not well-defined, due to the absence of a consistent definition of temperature and electrochemical potential.
The average current, given in Eq. (1), can be decomposed as an energy integral over a spectral current. The spectral current is given by
@@
Since the precision is a further performance quantifier of interest, in addition to the expectation values of the currents, their fluctuations will be considered. The fluctuations of a current $I^{z}$ is obtained from scattering theory by taking the expectation value of an auto-correlation function [Blanter2000Sep]
$\tilde{S}^{z}(t,\tau)=\langle\left\{\delta\hat{I}^{z}(t+\tau),\delta\hat{I}^{z}(t)\right\}\rangle,$ (5)
where $\delta\hat{I}^{z}(t)=\hat{I}^{z}(t)-I^{z}$. The zero-frequency noise, which is directly related to the variance of the current, is then found by integrating the autocorrelator over the times, giving the explicit expression [Blanter2000Sep],
$S^{z}$ $=$ $\int_{-\infty}^{\infty}d\varepsilon\,z^{2}(\varepsilon)\mathcal{D}(\varepsilon)(g(\varepsilon)g^{-}(\varepsilon)+f(\varepsilon)f^{-}(\varepsilon))$ (6)
@@
$s^{z}(\varepsilon)$ $\equiv$ $s^{z}_{\mathrm{thermal}}(\varepsilon)$ (7)
$=$ $z^{2}(\varepsilon)\mathcal{D}(\varepsilon)(g(\varepsilon)g^{-}(\varepsilon)+f(\varepsilon)f^{-}(\varepsilon)),$
@@ II.3 Performance quantifiers
Our objective is to optimize the performance of energy converters. Therefore, a crucial first step is to identify meaningful performance quantifiers. The first important aspect for the performance is naturally the actual desired output. This is quantified here as a generic current, $I^{x}$, see the definition in Eq. (1), with current factor $x(\varepsilon)$. For the matter of clarity and without loss of generality, we here always take the output current to be positive, $I^{x}\geq 0$. Consequently, optimization of the output current, which we present in Sec. III.1 always consists of maximizing $I^{x}$. Concretely, this means that if the desired output is negative following the definition in Eq. (1), one needs to set $x(\varepsilon)\to-z(\varepsilon)$ for the optimization procedure described below. This ensures that the total and spectral currents convey the same physical meaning.
@@
The strategy of how to define a meaningful, dimensionless efficiency becomes clear from the concrete example considered in Sec. IV, where we study the cooling power in the left contact. Since the right, resource contact is not thermal, the resource is typically not power or a heat current, as one would expect from a standard Peltier element or absorption refrigerator. Instead, motivated by the constraints put on energy conversion by the second law, we aim at characterizing the resource by an entropy current $I^{y}\rightarrow-I^{\Sigma,\mathrm{R}}$ (which is positive when entropy is produced in contact R). The constraints from the first law ensure energy and particle conservation, but cannot yield insight into the heat transfer process. In the thermal (left) contact, the entropy reduction is proportional to the desired cooling power, namely to the heat current leaving the contact $I^{x}\rightarrow I^{Q,\mathrm{L}}=TI^{\Sigma,\mathrm{L}}$. The increase in entropy, $I^{\Sigma,\mathrm{R}}>0$, implies that the right side acts as a resource (namely allowing by means of the second law that entropy is reduced elsewhere). Using the constraint by the second law (2c), we hence have $\eta=\frac{1}{T}\frac{I^{Q,\mathrm{L}}}{-I^{\Sigma,\mathrm{R}}}\leq 1$ with $c_{\eta}=1/T$. Alternatives for quantifying the resource that have been considered previously could be free-energy currents [Hajiloo2020Oct, Manzano2020Dec, Tesser2023] (with respect to the temperature of the thermal contact), $I^{y}\rightarrow I^{\varepsilon,\mathrm{R}}-TI^{\Sigma,\mathrm{R}}$.
@@
Furthermore, it is important in nanoscale conductors that the output of the energy-conversion process is precise. We hence aim at reducing the noise, where the performance quantifier is the noise $S^{x}$ of the current $I^{x}$ compared to the actual desired (average) output $I^{x}$. This can be quantified by the noise-to-signal ratio
$\mathrm{NSR}=c_{\mathrm{NSR}}\frac{S^{x}}{I^{x}}{\color[rgb]{0,0,0},}$ (9)
@@
where $c_{\mathrm{NSR}}$ is again a normalization factor for the dimensionality. For charge currents, this ratio is known as the Fano factor [Blanter2000Sep], where the constant takes the value $c_{\mathrm{NSR}}=1/2e$.
@@
It is well known from the output and the efficiency of standard engines that not both of these goals can be reached simultaneously, but that trade-offs need to be found. Here, we will consider one particular trade-off inspired by the thermodynamic uncertainty relation (TUR) [Barato2015Apr, Pietzonka2018May], which states that precision is bounded by entropy production. This TUR has been formulated in terms of transport currents [Brandner2018Mar, Kheradsoud2019Aug], where it takes the form
$A_{\mathrm{TUR}}^{x,\sigma}=\frac{S^{x}\sigma}{(I^{x})^{2}}\geq 2k_{\mathrm{B}},$ (10)
with $\sigma=-I^{\Sigma,L}-I^{\Sigma,R}$ being the total entropy production rate. Note that this bound is valid only for Markov processes and can be broken in quantum transport [Brandner2018Mar, Tesser2024May], in particular in situations where the transmission probability is close to 1 in some energy intervals.
We here consider a TUR-inspired factor
$A^{x,y}=c_{A}\frac{S^{x}I^{y}}{(I^{x})^{2}}{\color[rgb]{0,0,0},}$ (11)
as a performance quantifier, where we again introduce a normalization factor $c_{A}$. Our definition of Eq. (11) is a more general trade-off factor, where $I^{y}$ is not specified from the start. Note that this factor is not bounded either—due to quantum effects and due to the fact that $I^{y}$ is a generic current and not necessarily the total entropy production rate.